A snack food is made by mixing 5lb of popcorn that costs $.80 per pound with caramel that cost $2.40 per pound. How much caramel is needed to make a mixture that costs $1.40 per pound?
Answer by checkley77(12844)

5*.80+2.40x=1.40(5+x)
4+2.4x=7+1.4x
2.4x-1.4x=7-4
x=3 pounds of carmal is used.
Proof:
5*.80+2.40*3=1.40(5+3)
4+7.2=1.40*8
11.2=11.2